# MultiHateLoc: Towards Temporal Localisation of Multimodal Hate Content in Online Videos

### v1

@@ Abstract.
The rapid growth of video content on platforms such as TikTok and YouTube has intensified the spread of multimodal hate speech, where harmful cues emerge subtly and asynchronously across visual, acoustic, and textual streams. Existing research primarily focuses on video-level classification, leaving the practically crucial task of temporal localisation, identifying when hateful segments occur, largely unaddressed. This challenge is even more noticeable under weak supervision, where only video-level labels are available, and static fusion or classification-based architectures struggle to capture cross-modal and temporal dynamics. To address these challenges, we propose MultiHateLoc, the first framework designed for weakly-supervised multimodal hate localisation. MultiHateLoc incorporates (1) modality-aware temporal encoders to model heterogeneous sequential patterns, including a tailored text-based preprocessing module for feature enhancement; (2) dynamic cross-modal fusion to adaptively emphasise the most informative modality at each moment and a cross- modal contrastive alignment strategy to enhance multimodal feature consistency; (3) a modality-aware MIL objective to identify discriminative segments under video-level supervision. Despite relying solely on coarse labels, MultiHateLoc produces fine-grained, interpretable frame-level predictions. Experiments on HateMM and MultiHateClip show that our method achieves state-of-the-art performance in the localisation task.
@@
copyright: none
@@ 1. Introduction
Early hate speech research primarily focused on text-only detection, leveraging linguistic features or deep neural networks to classify short social media posts such as tweets or comments  (badjatiya2017deep; del2017hate). However, harmful content rapidly evolved into visual and multimodal forms. Memes, in particular, combine imagery and overlaid text to convey coded insults, satire, or hateful insinuations. This shift motivated multimodal hate meme detection (e.g., Facebook Hateful Memes Challenge (kiela2020hateful), Hate Speech Detection (gomez2020exploring)), demonstrating that cross-modal modelling is essential because hateful intent often emerges only when multiple modalities are considered together.
The focus of multimodal learning has expanded toward video-based hate content, where harmful cues unfold across visual frames, spoken language, background audio, and on-screen text (zhang2025dehate; yue2025multimodal). Early approaches typically rely on partial modalities, including vision-only abusive behaviour recognition (jiang2023abusive) and text–audio fusion for spoken toxicity analysis (alves2023audio). Building on HateMM (das2023hatemm), several recent works further explore multimodal hate video classification (yang2025revealing). A robust fusion study systematically compares modality combinations and feature encoders on HateMM and Hateful Memes (robustMMHate2025), CMFusion (CMFUsion2025) employs channel-wise and modality-wise fusion to strengthen tri-modal representation learning, and MoRe (lang2025more) leverages retrieval-augmented evidence and a mixture-of-experts design to enhance robustness in short videos, while MM-HSD (mmhsd2025) and MultiHateGNN (yue2025multimodal) introduce stronger multimodal backbones with cross-modal attention or graph neural networks for video-level prediction. ImpliHateVid (rehman2025impliHateVid) extends the problem to implicit hate in videos and also reports results on HateMM, and few-shot multimodal methods (ma2025fewshot) address label scarcity on the same dataset.
@@
To address this gap, we study a new research task: multimodal hate content localisation, which aims to identify the precise temporal segments in which hate occurs across modalities, as shown in Fig. 1 (c). Rather than requiring detailed frame-level annotations, we explore this problem under a weakly supervised setting, inspired by recent advances in weakly supervised video anomaly detection (WSVAD) (see Fig. 1 (b)). WSVAD methods (sultani2018real; tian2021weakly; wu2024weakly) are effective in learning temporal patterns from video-level labels by aligning visual and audio cues over time. However, they typically overlook the text modality, which is critical in the context of multimodal hate detection. Therefore, a major challenge in our work is to extend weakly supervised approaches to explicitly incorporate temporal modelling of text features, enabling a more complete and fine-grained understanding of multimodal hate content. A second major challenge in localising multimodal hate content lies in the high variability of modality relevance across time. Hate cues may not be evenly distributed; text might be dominant in one segment, while audio or visual signals become more salient elsewhere. This context-dependent salience makes it ineffective to apply static or early fusion strategies.
These challenges motivate our design of MultiHateLoc, as shown in Fig 2, the first multimodal hate localisation framework which incorporates 1) Modality-aware temporal modeling to capture intra-modal sequential information across dense video frames, audio features, and text sequences. We further proposed a novel text embedding module for sentence-wise text feature extraction. 2) Dynamic cross-modal fusion to adaptively emphasise the modality contributions based on context at each time-step and a cross-modal contrastive alignment strategy to enhance multimodal feature consistency. 3) Modality-aware top-K Multiple Instance Learning (MIL) objective that guides weak supervision toward the most relevant segments. Together, these components overcome the inherent limitations of prior work and enable fine-grained localisation of multimodal hate content. We evaluate MultiHateLoc on the HateMM (das2023hatemm) and MultiHateClip (wang2024multihateclip) datasets, achieving state-of-the-art localisation performance. To the best of our knowledge, this is the first work to provide a fully tri-modal, weakly-supervised, and temporally precise solution for hate content localisation in online videos.
@@ 2.1. Hate Speech Detection
Early research on hate speech detection has largely centered around textual content, where a wide array of machine learning and deep learning techniques have been explored. For example, MacAvaney et al. (macavaney2019hate) employed a multi-view SVM approach that combines diverse feature representations to improve both accuracy and interpretability. Badjatiya et al. (badjatiya2017deep) investigated a range of neural models, including FastText, CNNs, and LSTMs, demonstrating their effectiveness on social media datasets such as tweets. Other conventional techniques, such as support vector machines, have also been utilized for hate speech classification (del2017hate; sanoussi2022detection).
Despite promising results, these approaches are fundamentally limited by their reliance on a single modality — text. In real-world online discourse, hate speech can be highly nuanced, leveraging sarcasm, coded language, or context-dependent cues that are difficult to capture with text alone. This makes purely textual models susceptible to false negatives, particularly when hateful intent is implied rather than explicit (yin2021towards; rottger2020hatecheck; grondahl2018all). These limitations underscore the need for models capable of integrating richer contextual and multimodal signals to improve robustness and generalizability.
@@ 2.2. Multimodal Hateful Content Detection
Multimodal approaches substantially outperform unimodal systems for hateful content detection. Early multimodal works combined image–text or video–text signals using concatenation, gating, or bilinear pooling (yang2019exploring; gomez2020exploring; chhabra2023multimodal), and transformer-based architectures such as ViLBERT further improved cross-modal alignment (kiela2020hateful). More recent efforts extend to tri-modal settings, with HateMM (das2023hatemm) demonstrating the benefits of jointly modeling video, audio, and text. Two main directions have emerged: (1) adapting pre-trained multimodal models, e.g., CLIP-based systems like VAD-CLIP (wu2023vadclip), which incorporate temporal modeling but still rely on global text features; and (2) task-specific fusion architectures such as MM-HSD (cespedes2025mm), CMFusion (CMFUsion2025), MoRE (lang2025biting) MultiHateGNN (yue2025multimodal). More recent works such as (yang2025revealing) explore the label noise in hate videos and in DeHate(zhang2025dehate), a novel fine-grained large-scale hate content dataset has been proposed for advanced hate detection. While these methods advance video-level classification, a recent survey (hee2024recent) highlights persistent temporal misalignment across modalities and the lack of fine-grained grounding, especially for text and audio cues. Consequently, existing multimodal hate detection methods remain classification-centric, leaving the problem of temporal localisation largely unexplored.
@@ 2.3. Weakly Supervised Video Anomaly Detection
Frame-level annotations are costly to obtain, so weakly supervised learning has become a common solution for temporal localisation tasks. Multi-Instance Learning (MIL) is widely used in this setting. It trains models with only video-level labels and allows them to infer the segments that contain target events. MIL-based methods have shown strong performance in action recognition and video anomaly detection. Examples include the MIL ranking framework by Sultani et al. (sultani2018real), self-supervised anomaly modeling by Rudolph et al. (Rudolph_2021_WACV), and GNN-based approaches for group-level anomalies (ullah2023ad; li2022weakly). These works demonstrate that weak supervision can capture temporal patterns without detailed annotations. However, most research focuses on explicit visual anomalies in surveillance videos, which differs from the subtle and often cross-modal nature of hateful content.
Hate signals are frequently concealed. A video may look visually normal but contain harmful cues in the audio or text. Existing WS-VAD methods (sultani2018real; li2022weakly; wu2023vadclip) rely mainly on visual features, so they struggle to capture such multimodal patterns. To address this limitation, we propose a weakly supervised MIL framework that integrates video, audio, and textual features. It uses only video-level labels but can still predict the temporal boundaries of hateful segments. This approach enables fine-grained localisation under weak supervision and fills an important gap in applying WS-VAD techniques to multimodal hate content.
@@ 3.1. Problem Formulation
  • Video ($F_{v}\in\mathbb{R}^{T\times 768}$): We extract frame-level visual features using a pre-trained ViT-B/16 (dosovitskiy2020image). Each frame is processed independently to retain spatial details.
  • Audio ($F_{a}\in\mathbb{R}^{T\times 128}$): Audio features are extracted from 1-second clips using VGGish (hochreiter1997long). The resulting sequence is linearly interpolated to match the video length $T$. This produces a temporally aligned audio representation with the same frame-level resolution as the visual stream.
  • Text ($F_{l}\in\mathbb{R}^{T\times 768}$): We propose a novel sentence-wise text embeddings for text-based localisation, as shown in Fig. 3, which generated via four steps: (1) Transcribe video audio to text using Whisper (radford2023robust) ; (2) Split transcribed text into $S$ sentence-wise fragments based on each sentence start and end timestamps; (3) Encode each sentence with BERT (devlin2018bert) to obtain a 768-dimensional feature; (4) Expand sentence-level features to the frame length $T$ by repeated padding within each sentence’s timestamp interval. This converts $S$ sentence embeddings into frame-aligned text features and resolves temporal misalignment in naive text encodings.
@@ 4. Experiments
We utilise two multimodal datasets: HateMM(das2023hatemm) and MultiHateClip(wang2024multihateclip), both of which contain video, audio, and text. HateMM includes  43 hours of videos from BitChute, labeled as “hateful” or “non-hateful.” For hateful videos, frame-level annotations mark specific segments. The dataset contains 431 hateful and 652 non-hateful videos, supporting both classification and localization tasks. MultiHateClip contains 2,000 videos from YouTube and Bilibili, evenly divided into hateful, offensive, and normal categories. Each video has segment-level labels and additional metadata (e.g., target group, modality, audience). It supports multilingual and fine-grained hate detection.
@@ 4.2. Baseline Comparisons
Weakly-supervised anomaly localisation (VAD-CLIP (wu2023vadclip)): These models naturally output temporal saliency curves under video-level labels, making them the closest analogue to our weakly-supervised setup. However, they lack the explicit cross-modal modelling and semantic alignment required for hate content.
@@
Modern multimodal classifier (CMFusion (CMFUsion2025)): We include CMFusion to test whether a state-of-the-art tri-modal classifier can be adapted into a temporal localiser via a sliding-window strategy. This baseline allows us to contrast classification-driven fusion with our temporal, modality-aware approach.
@@
Additionally, the comparison between modality combinations (e.g., V+A, V+T, T+A) reveals that MultiHateLoc effectively leverages complementary information among modalities, achieving consistently better results than the single or dual modality configurations in VAD-CLIP. This underscores the importance and effectiveness of our dynamic modality selection and cross-modal fusion strategies. We also compare our model with up-to-data classification model CMFusion (CMFUsion2025), regarding to the fusion part, our model also consistently achieves better performance.
@@ 4.3.2. Text Encoding Strategy
This section verifies the effectiveness of our sentence-wise text encoding by comparing it with naive text processing (common in prior works (das2023hatemm; wang2024multihateclip) and text-omitted settings.
@@
VAD-CLIP(wu2023vadclip) + W/O Text 0.546 0.752
VAD-CLIP(wu2023vadclip) + Naive Text 0.551 0.770
VAD-CLIP(wu2023vadclip) + Sentence-wise Text 0.578 0.779
@@ 4.3.4. Module Simplification
Simplification of Temporal Modeling
Table 6. Comparison of different temporal encoders used in our framework.
Temporal Encoder AP AUC
RNN-based Temporal Model 0.612 0.770
CNN-based Temporal Model 0.600 0.762
Transformer-based 0.645 0.799
We also investigate the impact of simplifying the Transformer-based temporal encoder. Table 6 reports the results of replacing the Transformer with simpler RNN or CNN-based temporal models. The Transformer achieves the best performance with an AP of 0.635, while RNN and CNN-based alternatives yield lower AP scores of 0.612 and 0.600, respectively. This confirms the superiority of Transformer-based temporal encoding for our task.

### v2

@@ Abstract.
The rapid growth of video content on platforms such as TikTok and YouTube has intensified the spread of multimodal hate speech, where harmful cues emerge subtly and asynchronously across visual, acoustic, and textual streams. Existing research primarily focuses on video-level classification, leaving the practically crucial task of temporal localisation, identifying when hateful segments occur, largely unaddressed. This challenge is even more noticeable under weak supervision, where only video-level labels are available, and static fusion or classification-based architectures struggle to capture cross-modal and temporal dynamics. To address these challenges, we propose MultiHateLoc, the first framework designed for weakly-supervised multimodal hate localisation. MultiHateLoc incorporates (1) modality-aware temporal encoders to model heterogeneous sequential patterns, including a tailored text-based preprocessing module for feature enhancement; (2) dynamic cross-modal fusion to adaptively emphasise the most informative modality at each moment and a cross- modal contrastive alignment strategy to enhance multimodal feature consistency; (3) a modality-aware MIL objective to identify discriminative segments under video-level supervision. Despite relying solely on coarse labels, MultiHateLoc produces fine-grained, interpretable frame-level predictions. Experiments on HateMM and MultiHateClip show that our method achieves state-of-the-art performance in the localisation task. Code is available at https://github.com/mmilabuk/multihateloc.
@@
Multimodal; Hateful Content Detection; Temporal Localisation
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3793032isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Speech recognitionccs: Computing methodologies Scene anomaly detection
@@ 1. Introduction
Early hate speech research primarily focused on text-only detection, leveraging linguistic features or deep neural networks to classify short social media posts such as tweets or comments  (Badjatiya et al., 2017; Del Vigna12 et al., 2017). However, harmful content rapidly evolved into visual and multimodal forms. Memes combine imagery and overlaid text to convey coded insults, satire, or hateful insinuations. This shift motivated multimodal hate meme detection (e.g., Facebook Hateful Memes Challenge (Kiela et al., 2020), Hate Speech Detection (Gomez et al., 2020)), demonstrating that cross-modal modelling is essential because hateful intent often emerges only when multiple modalities are considered together.
The focus of multimodal learning has expanded toward video-based hate content, where harmful cues unfold across visual frames, spoken language, background audio, and on-screen text (Zhang et al., 2025; Yue et al., 2025). Early approaches typically rely on partial modalities, including vision-only abusive behaviour recognition (Jiang et al., 2023) and text–audio fusion for spoken toxicity analysis (Alves et al., 2023). Building on HateMM (Das et al., 2023), several recent works further explore multimodal hate video classification (Yang et al., 2025). A robust fusion study systematically compares modality combinations and feature encoders on HateMM and Hateful Memes (Author and others, 2025), CMFusion (Zhang et al., 2024) employs channel-wise and modality-wise fusion to strengthen tri-modal representation learning, and MoRe (Lang et al., 2025a) leverages retrieval-augmented evidence and a mixture-of-experts design to enhance robustness in short videos, while MM-HSD (Céspedes-Sarrias et al., 2025b) and MultiHateGNN (Yue et al., 2025) introduce stronger multimodal backbones with cross-modal attention or graph neural networks for video-level prediction. ImpliHateVid (Rehman et al., 2025) extends to implicit hate in videos and also reports results on HateMM, and few-shot multimodal methods (Ma and others, 2025) address label scarcity on the same dataset.
@@
To address this gap, we study a new research task: multimodal hate content localisation, which aims to identify the precise temporal segments in which hate occurs across modalities, as shown in Fig. 1 (c). Rather than requiring detailed frame-level annotations, we explore this problem under a weakly supervised setting, inspired by recent advances in weakly supervised video anomaly detection (WSVAD) (see Fig. 1 (b)). WSVAD methods (Sultani et al., 2018; Tian et al., 2021; Wu et al., 2024a) are effective in learning temporal patterns from video-level labels by aligning visual and audio cues over time. However, they typically overlook the text modality, which is critical in the context of multimodal hate detection. Therefore, a major challenge in our work is to extend weakly supervised approaches to explicitly incorporate temporal modelling of text features, enabling a more complete and fine-grained understanding of multimodal hate content. A second major challenge in localising multimodal hate content lies in the high variability of modality relevance across time. Hate cues may not be evenly distributed; text might be dominant in one segment, while audio or visual signals become more salient elsewhere. This context-dependent salience makes it ineffective to apply static or early fusion strategies.
These challenges motivate our design of MultiHateLoc, as shown in Fig 2, the first multimodal hate localisation framework which incorporates 1) Modality-aware temporal modeling to capture intra-modal sequential information across dense video frames, audio features, and text sequences. We further proposed a novel text embedding module for sentence-wise text feature extraction. 2) Dynamic cross-modal fusion to adaptively emphasise the modality contributions based on context at each time-step and a cross-modal contrastive alignment strategy to enhance multimodal feature consistency. 3) Modality-aware top-K Multiple Instance Learning (MIL) objective that guides weak supervision toward the most relevant segments. Together, these components overcome the inherent limitations of prior work and enable fine-grained localisation of multimodal hate content. We evaluate MultiHateLoc on the HateMM (Das et al., 2023) and MultiHateClip (Wang et al., 2024) datasets, achieving state-of-the-art localisation performance. To the best of our knowledge, this is the first work to provide a fully tri-modal, weakly-supervised, and temporally precise solution for hate content localisation in online videos.
@@ 2.1. Hate Speech Detection
Early research on hate speech detection has largely centered around textual content, where a wide array of machine learning and deep learning techniques have been explored. For example, MacAvaney et al. (MacAvaney et al., 2019) employed a multi-view SVM approach that combines diverse feature representations to improve both accuracy and interpretability. Badjatiya et al. (Badjatiya et al., 2017) investigated a range of neural models, including FastText, CNNs, and LSTMs, demonstrating their effectiveness on social media datasets such as tweets. Other conventional techniques, such as support vector machines, have also been utilized for hate speech classification (Del Vigna12 et al., 2017; Sanoussi et al., 2022).
Despite promising results, these approaches are fundamentally limited by their reliance on a single modality — text. In real-world online discourse, hate speech can be highly nuanced, leveraging sarcasm, coded language, or context-dependent cues that are difficult to capture with text alone. This makes purely textual models susceptible to false negatives, particularly when hateful intent is implied rather than explicit (Yin and Zubiaga, 2021; Röttger et al., 2020; Gröndahl et al., 2018). These limitations underscore the need for models capable of integrating richer contextual and multimodal signals to improve robustness and generalizability.
@@ 2.2. Multimodal Hateful Content Detection
Multimodal approaches substantially outperform unimodal systems for hateful content detection. Early multimodal works combined image–text or video–text signals using concatenation, gating, or bilinear pooling (Yang et al., 2019; Gomez et al., 2020; Chhabra and Vishwakarma, 2023), and transformer-based architectures such as ViLBERT further improved cross-modal alignment (Kiela et al., 2020). More recent efforts extend to tri-modal settings, with HateMM (Das et al., 2023) demonstrating the benefits of jointly modeling video, audio, and text. Two main directions have emerged: (1) adapting pre-trained multimodal models, e.g., CLIP-based systems like VAD-CLIP (Wu et al., 2024b), which incorporate temporal modeling but still rely on global text features; and (2) task-specific fusion architectures such as MM-HSD (Céspedes-Sarrias et al., 2025a), CMFusion (Zhang et al., 2024), MoRE (Lang et al., 2025b) MultiHateGNN (Yue et al., 2025). More recent works such as (Yang et al., 2025) explore the label noise in hate videos and in DeHate(Zhang et al., 2025), a novel fine-grained large-scale hate content dataset has been proposed for advanced hate detection. While these methods advance video-level classification, a recent survey (Koushik et al., 2025) highlights persistent temporal misalignment across modalities and the lack of fine-grained grounding, especially for text and audio cues. Consequently, existing multimodal hate detection methods remain classification-centric, leaving the problem of temporal localisation largely unexplored.
@@ 2.3. Weakly Supervised Video Anomaly Detection
Frame-level annotations are costly to obtain, so weakly supervised learning has become a common solution for temporal localisation tasks. Multi-Instance Learning (MIL) is widely used in this setting. It trains models with only video-level labels and allows them to infer the segments that contain target events. MIL-based methods have shown strong performance in action recognition and video anomaly detection. Examples include the MIL ranking framework by Sultani et al. (Sultani et al., 2018), self-supervised anomaly modeling by Rudolph et al. (Rudolph et al., 2021), and GNN-based approaches for group-level anomalies (Ullah et al., 2023; Li et al., 2022). These works demonstrate that weak supervision can capture temporal patterns without detailed annotations. However, most research focuses on explicit visual anomalies in surveillance videos, which differs from the subtle and often cross-modal nature of hateful content.
Hate signals are frequently concealed. A video may look visually normal but contain harmful cues in the audio or text. Existing WS-VAD methods (Sultani et al., 2018; Li et al., 2022; Wu et al., 2024b) rely mainly on visual features, so they struggle to capture such multimodal patterns. To address this limitation, we propose a weakly supervised MIL framework that integrates video, audio, and textual features. It uses only video-level labels but can still predict the temporal boundaries of hateful segments. This approach enables fine-grained localisation under weak supervision and fills an important gap in applying WS-VAD techniques to multimodal hate content.
@@ 3.1. Problem Formulation
  • Video ($F_{v}\in\mathbb{R}^{T\times 768}$): We extract frame-level visual features using a pre-trained ViT-B/16 (Dosovitskiy et al., 2020). Each frame is processed independently to retain spatial details.
  • Audio ($F_{a}\in\mathbb{R}^{T\times 128}$): Audio features are extracted from 1-second clips using VGGish (Hochreiter, 1997). The resulting sequence is linearly interpolated to match the video length $T$. This produces a temporally aligned audio representation with the same frame-level resolution as the visual stream.
  • Text ($F_{l}\in\mathbb{R}^{T\times 768}$): We propose a novel sentence-wise text embeddings for text-based localisation, as shown in Fig. 3, which generated via four steps: (1) Transcribe video audio to text using Whisper (Radford et al., 2023) ; (2) Split transcribed text into $S$ sentence-wise fragments based on each sentence start and end timestamps; (3) Encode each sentence with BERT (Devlin et al., 2018) to obtain a 768-dimensional feature; (4) Expand sentence-level features to the frame length $T$ by repeated padding within each sentence’s timestamp interval. This converts $S$ sentence embeddings into frame-aligned text features and resolves temporal misalignment in naive text encodings.
@@ 4. Experiments
We utilise two multimodal datasets: HateMM(Das et al., 2023) and MultiHateClip(Wang et al., 2024), both of which contain video, audio, and text. HateMM includes  43 hours of videos from BitChute, labeled as “hateful” or “non-hateful.” For hateful videos, frame-level annotations mark specific segments. The dataset contains 431 hateful and 652 non-hateful videos, supporting both classification and localization tasks. MultiHateClip contains 2,000 videos from YouTube and Bilibili, evenly divided into hateful, offensive, and normal categories. Each video has segment-level labels and additional metadata (e.g., target group, modality, audience). It supports multilingual and fine-grained hate detection.
@@ 4.2. Baseline Comparisons
Weakly-supervised anomaly localisation (VAD-CLIP (Wu et al., 2024b)): These models naturally output temporal saliency curves under video-level labels, making them the closest analogue to our weakly-supervised setup. However, they lack the explicit cross-modal modelling and semantic alignment required for hate content.
@@
Modern multimodal classifier (CMFusion (Zhang et al., 2024)): We include CMFusion to test whether a state-of-the-art tri-modal classifier can be adapted into a temporal localiser via a sliding-window strategy. This baseline allows us to contrast classification-driven fusion with our temporal, modality-aware approach.
@@
Additionally, the comparison between modality combinations (e.g., V+A, V+T, T+A) reveals that MultiHateLoc effectively leverages complementary information among modalities, achieving consistently better results than the single or dual modality configurations in VAD-CLIP. This underscores the importance and effectiveness of our dynamic modality selection and cross-modal fusion strategies. We also compare our model with up-to-data classification model CMFusion (Zhang et al., 2024), regarding to the fusion part, our model also consistently achieves better performance.
@@ 4.3.2. Text Encoding Strategy
This section verifies the effectiveness of our sentence-wise text encoding by comparing it with naive text processing (common in prior works (Das et al., 2023; Wang et al., 2024) and text-omitted settings.
@@
VAD-CLIP(Wu et al., 2024b) + W/O Text 0.546 0.752
VAD-CLIP(Wu et al., 2024b) + Naive Text 0.551 0.770
VAD-CLIP(Wu et al., 2024b) + Sentence-wise Text 0.578 0.779